The general manager of Frozenaire Refrigerator Company determines that on Monday it cost the company a total of $25,000 to build 30 refrigerators and on Tuesday it cost $30,000 to build 40 refrigerators. Determine for him the company's daily fixed costs and their marginal cost (that is, the cost to build one refrigerator).
Answer by foleysdad(7)

We are trying to find two answers here. The cost per refrigerator and the overhead cost of the factory per day.
Let's assume the following variables:
r = cost per refrigerator
f = fixed costs (or overhead) per day for the factory
Monday is 30r + f = 25000
Tuesday is 40r + f = 30000
Lets solve for f using Monday
f= 25000 -30r
Now that we know what f is in terms of r let's put it into Tuesday's equation.
40r + (25000 - 30r) = 30000
40r -30r + 25000 = 30000
10r + 25000 = 30000
10r = 5000
r = 500
Plug r into either equation.
30(500) + f = 25000
15000 + f = 25000
f = 10000
or
40(500) + f = 30000
20000 + f = 30000
f = 10000
The answers are :
f (overhead) = 10,000 per day
r (cost per refrigerator) = 500 per unit